OpenStudy (anonymous):

Hi can someone algebraically walk me through this LIMIT problem WITHOUT using l'Hospital's rule: lim(1/(x sqrt(1+x)) - (1/x)) as x -> 0

OpenStudy (anonymous):

\[\frac{1}{x\sqrt{1+x}}-\frac{1}{x}\]?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

hmmm first think i would do is subtract

OpenStudy (anonymous):

ya i tried that, then it still looks really wierd

OpenStudy (anonymous):

did you get \[\frac{1-\sqrt{1+x}}{x\sqrt{1+x}}\]?

OpenStudy (anonymous):

ya on my paper

OpenStudy (anonymous):

that's the confusing part for me. .. . i'm not sure what to do after

OpenStudy (anonymous):

because usually i just factor or divide into something perfectly

OpenStudy (anonymous):

not this time frequently you have to multiply by the conjugate to rationalize the numerator

OpenStudy (anonymous):

ohh is that what i'm suppose to do here?

OpenStudy (anonymous):

i didn't try that yet

OpenStudy (anonymous):

yes, multiply top and bottom by \(1+\sqrt{1+x}\) leave the denominator in factored form, don't multiply it out

OpenStudy (anonymous):

k thanks again satellite i think i got it now!

OpenStudy (anonymous):

yw